# Primal-Dual iLQR for GPU-Accelerated Learning and Control in Legged Robots

### v2

@@ II-D Parallel Associative Scan
for $i=1,\dots,N$ while for $i=0$ we set $\boldsymbol{\bar{A}}_{0,1}=\boldsymbol{0}$ and $\boldsymbol{\bar{b}}_{0,1}=\boldsymbol{\bar{A}}_{0}\boldsymbol{\delta x}_{0}+\boldsymbol{\bar{b}}_{0}$. From the parallel scan, we obtained the updated optimal state as:
@@ II-E Parallel Line Search
Once we update the optimal directions $\boldsymbol{\delta x}$, $\boldsymbol{\delta u}$, and $\boldsymbol{\delta\lambda}$, we can perform a backtracking line search to calculate the step length $\alpha$ to finally get the new guess as:
$\boldsymbol{x}^{k+1}=\boldsymbol{x}^{k}+\alpha\boldsymbol{\delta x}$ $\boldsymbol{\lambda}^{k+1}=\boldsymbol{\lambda}^{k}+\alpha\boldsymbol{\delta\lambda}$ (18)
$\boldsymbol{u}^{k+1}=\boldsymbol{u}^{k}+\alpha\boldsymbol{\delta u}$
The SQP linear rollout allows for the computation to be easily performed in parallel for the horizon. To evaluate $\alpha$, we implemented a filter line search similar to the one implemented in [ipopt]. The filter line search ensures that the step taken in the iteration reduces the constraint violation or the cost. Such a method has already been successfully deployed for legged robot control in [Perceptive_based_MPC]. We measure the satisfaction of the dynamics constraints as:
$\theta=\sum_{i=0}^{N+1}\|\boldsymbol{x}_{i+1}-\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})\|_{2}$ (19)
@@ II-F Practical Implementation
$\boldsymbol{Q}_{N+1}$ $=\nabla_{x}\boldsymbol{\epsilon}_{N+1}^{\top}W_{T}\nabla_{x}\boldsymbol{\epsilon}_{N+1}$ $\boldsymbol{S}_{i}=\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{u}\boldsymbol{\epsilon}_{i}$ (20)
@@
$\mathcal{B}({\color[rgb]{0,0,0}\xi},\mu,\delta)\begin{cases}-\mu\ln({\color[rgb]{0,0,0}\xi}),&\text{if }{\color[rgb]{0,0,0}\xi}\geq\delta,\\ \frac{\mu}{2}\left(\left(\frac{{\color[rgb]{0,0,0}\xi}-2\delta}{\delta}\right)^{2}-1\right)-\mu\ln(\delta),&\text{if }{\color[rgb]{0,0,0}\xi}<\delta.\end{cases}$ (21)
@@
$\boldsymbol{Q}_{N+1}$ $={\color[rgb]{0,0,0}\nabla_{x}\boldsymbol{\epsilon}_{N+1}^{\top}W_{T}\nabla_{x}\boldsymbol{\epsilon}_{N+1}}+\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{N+1}^{\top}\nabla_{{\color[rgb]{0,0,0}\xi}{\color[rgb]{0,0,0}\xi}}\mathcal{B}_{T}\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{N+1}$ (22)
@@ IV-A Performance evaluation
To finally validate the MPC, we also performed experiments on the real hardware, as shown in Fig. 1 and in the accompanying video. For the real experiments, the solver runs on an Nvidia RTX 4050 laptop GPU. Figure 1 shows snapshots of the robot walking blindly up the steps and through uneven terrain using only proprioceptive data for state estimation [muse]. We also evaluated the proposed approach when used to control two different robots at the same time in a centralized control implementation in simulation. In this example, the two robots are controlled by the same MPC that evaluates the control input for both systems at 5050Hz. The MPC also includes a collision avoidance constraint, in the form of a quadratic penalty term. In the tested scenario, the two robots are tasked to track two perpendicular trajectories that cross each other and would lead to a collision. Figure 3 shows the resulting path of the two robots, where the dashed grey lines represent the distance between the two robots. As shown, both robots deviate from their desired paths to avoid collisions, demonstrating coordinated behavior. While being just a simple example, this result highlights the potential of our MPC in solving collaborative tasks, further implementation in more complex scenarios, such as collaborative carrying is left for future work. Figure 4 presents the average solving time breakdown for the three solvers in the quadruped trotting scenario with a horizon of N=100. It highlights the sequential backward pass in crocoddyl, the QP solver call in acados, and the parallel scan in our method for evaluating the search direction. A minor computational overhead of approximately $1ms$ arises from GPU communication, as only the initial state is sent and the commanded torques, joint positions, and velocities are retrieved.

### v3

@@ II-D Parallel Associative Scan
for $i=1,\dots,N$, while for $i=0$ we set $\boldsymbol{\bar{A}}_{0,1}=\boldsymbol{0}$ and $\boldsymbol{\bar{b}}_{0,1}=\boldsymbol{\bar{A}}_{0}\boldsymbol{\delta x}_{0}+\boldsymbol{\bar{b}}_{0}$. From the parallel scan, we obtained the updated optimal state as:
@@ II-E Parallel Line Search
Once we update the optimal directions $\boldsymbol{\delta x}$, $\boldsymbol{\delta u}$, and $\boldsymbol{\delta\lambda}$, we can perform a backtracking line search to calculate the step length $\alpha$ to finally get the new guess as
$\boldsymbol{x}^{k+1}=\boldsymbol{x}^{k}+\alpha\boldsymbol{\delta x},$ $\boldsymbol{\lambda}^{k+1}=\boldsymbol{\lambda}^{k}+\alpha\boldsymbol{\delta\lambda},$ (18)
$\boldsymbol{u}^{k+1}=\boldsymbol{u}^{k}+\alpha\boldsymbol{\delta u}.$
The SQP linear rollout allows for the computation to be easily performed in parallel for the horizon. To evaluate $\alpha$, we implemented a filter line search similar to the one implemented in [ipopt]. The filter line search ensures that the step taken in the iteration reduces the constraint violation or the cost. Such a method has already been successfully deployed for legged robot control in [Perceptive_based_MPC]. We measure the satisfaction of the dynamics constraints as
$\theta=\sum_{i=0}^{N+1}\|\boldsymbol{x}_{i+1}-\boldsymbol{h}(\boldsymbol{x}_{i},\boldsymbol{u}_{i})\|_{2}.$ (19)
@@ II-F Practical Implementation
$\boldsymbol{Q}_{N+1}$ $=\nabla_{x}\boldsymbol{\epsilon}_{N+1}^{\top}W_{N+1}\nabla_{x}\boldsymbol{\epsilon}_{N+1}$ $\boldsymbol{S}_{i}=\nabla_{x}\boldsymbol{\epsilon}_{i}^{\top}W_{i}\nabla_{u}\boldsymbol{\epsilon}_{i}$ (20)
@@
$\mathcal{B}({\color[rgb]{0,0,0}\xi},\mu,\delta)\begin{cases}-\mu\ln({\color[rgb]{0,0,0}\xi}),&\text{if }{\color[rgb]{0,0,0}\xi}\geq\delta,\\ \frac{\mu}{2}\left(\left(\frac{{\color[rgb]{0,0,0}\xi}-2\delta}{\delta}\right)^{2}-1\right)-\mu\ln(\delta),&\text{if }{\color[rgb]{0,0,0}\xi}<\delta,\end{cases}$ (21)
@@
$\boldsymbol{Q}_{N+1}$ $={\color[rgb]{0,0,0}\nabla_{x}\boldsymbol{\epsilon}_{N+1}^{\top}W_{N+1}\nabla_{x}\boldsymbol{\epsilon}_{N+1}}+\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{N+1}^{\top}\nabla_{{\color[rgb]{0,0,0}\xi}{\color[rgb]{0,0,0}\xi}}\mathcal{B}_{N+1}\nabla_{x}\boldsymbol{{\color[rgb]{0,0,0}\xi}}_{N+1}$ (22)
@@ IV-A Performance evaluation
To finally validate the MPC, we also performed experiments on the real hardware, as shown in Fig. 1 and in the accompanying video. For the real experiments, the solver runs on an Nvidia RTX 4050 laptop GPU. Figure 1 shows snapshots of the robot walking blindly up the steps and through uneven terrain using only proprioceptive data for state estimation [muse]. We also evaluated the proposed approach when used to control two different robots at the same time in a centralized control implementation in simulation. In this example, the two robots are controlled by the same MPC that evaluates the control input for both systems at 5050Hz. The MPC also includes a collision avoidance constraint, in the form of a quadratic penalty term. In the tested scenario, the two robots are tasked to track two perpendicular trajectories that cross each other and would lead to a collision. Figure 3 shows the resulting path of the two robots, where the dashed grey lines represent the distance between the two robots. As shown, both robots deviate from their desired paths to avoid collisions, demonstrating coordinated behavior. While being just a simple example, this result highlights the potential of our MPC in solving collaborative tasks. Further implementation in more complex scenarios, such as collaborative carrying, is left for future work. Figure 4 presents the average solving time breakdown for the three solvers in the quadruped trotting scenario with a horizon of N=100. It highlights the sequential backward pass in crocoddyl, the QP solver call in acados, and the parallel scan in our method for evaluating the search direction. A minor computational overhead of approximately $1ms$ arises from GPU communication, as only the initial state is sent and the commanded torques, joint positions, and velocities are retrieved.